# bibliography.bib

### v1

@@ Abstract
We draw from Raney duality and generalize the notion of canonical extension for distributive lattices to the context of frames. This is a generalization because for distributive lattices we are in the context of coherent spaces, whereas frames represent arbitrary spaces. We introduce Raney extensions of frames, pairs $(L,C)$ where $C$ is a coframe, $L\subseteq C$ is a frame that meet-generates it, and the inclusion $L\subseteq C$ preserves the frame operations as well as the strongly exact meets. We think of these as being algebraic versions of the embedding $\Omega(X)\subseteq\mathcal{U}(X)$ of a frame of opens into a lattice of saturated sets (upper sets in the specialization order). We show that a frame may have several Raney extensions, and all satisfy a generalization of the properties of density and compactness from the theory of canonical extensions. We show that every frame $L$ has the largest and the smallest Raney extension, and that these are, respectively, the coframe of fitted sublocales $\mathsf{S}_{\mathfrak{o}}(L)$ and the opposite of the frame $\mathsf{S}_{\mathfrak{c}}(L)$ of joins of closed sublocales (both studied in [moshier20] and [moshier22]). We thus show that these structures have universal properties which are dual of one another. For Raney extensions $(L,C)$ and $(M,D)$, we characterize the frame morphisms $f:L\to M$ which can be extended to a morphism of Raney extensions. We apply this result to obtain a characterization of morphisms of frames $f:L\to M$ which can be lifted to frame morphisms $\mathsf{S}_{\mathfrak{c}}(L)\to\mathsf{S}_{\mathfrak{c}}(M)$. We also show that the canonical extension of a locally compact frame, as defined in [jakl20], is the free Raney extension over it such that it is algebraic (generated by its compact elements). We give a characterization of sobriety and of strict sobriety based on a variation of the compactness property of the extension $(\Omega(X),\mathcal{U}(X))$.
@@ 1 Introduction
In this work, we study an algebraic, pointfree version of the embedding of a frame of open sets into a lattice of saturated sets. Our constructions are inspired by Raney duality, as illustrated in [bezhanishvili20]. In Raney duality, rather than mapping a space $X$ to the frame $\Omega(X)$ of its open sets as one does in pointfree topology, we map it to the embedding $\Omega(X)\subseteq\mathcal{U}(X)$ of its open sets into the lattice of saturated sets [^1] Saturated sets are intersections of open sets.. In this duality, on the algebraic side all objects of the category have, so to speak, enough points. That is, every Raney algebra is of the form $(\Omega(X),\mathcal{U}(X))$ for some space $X$. In this work, we extend the category of Raney algebras to include pointfree objects. We also build on the notion of canonical extension of a distributive lattice: an algebraic, pointfree version of the embedding of the lattice of compact open sets of a coherent space into the lattice of saturated sets. For a distributive lattice, its canonical extension is unique. For frames, we claim, there is no unique way of extending a frame $L$ to a pointfree lattice of saturated sets.
@@
We will consider as objects of our category Raney extensions, pairs $(L,C)$ where $C$ is a coframe and $L\subseteq C$ is a frame which meet-generates $C$ and such that the embedding preserves the frame operations together with strongly exact meets [^2] Strongly exact meets are the pointfree version of those intersections of open sets which are open. Because a meet of a collection $\{U_{i}:i\in I\}$ of opens in general is calculated as the interior of $\bigcap_{i}U_{i}$, these are exactly the meets that are preserved by the embedding $\Omega(X)\subseteq\mathcal{U}(X)$.. We use classical results from the theory of polarities of Birkhoff to show that all Raney extensions are, up to isomorphism, pairs of the form $(L,\mathcal{F}^{op})$, where $\mathcal{F}\subseteq\mathsf{Filt}(L)$ is a sublocale of the collection of filters of $L$. We build on results in [jakl24] to show that all Raney extensions have universal properties which are variations of the universal property of the canonical extension, by considering variations of the density and compactness property. In particular, for a collection of filters $\mathcal{F}\subseteq L$ of a frame $L$ we will say that a Raney extension $(L,C)$ is
  • dense with respect to $\mathcal{F}$ if every element of $C$ is a join of elements of the form $\bigwedge F$ for $F\in\mathcal{F}$;
  • compact with respect to $\mathcal{F}$ if for each $F\in\mathcal{F}$ and each $a\in L$ we have that $\bigwedge F\leq a$ implies $a\in F$.
We also tackle the problem of lifting of morphisms. For Raney extensions $(L,C)$ and $(M,D)$, we will characterize frame morphisms $f:L\to M$ such that can be extended to morphisms of Raney extensions. Recently, the collection $\mathsf{S}_{\mathfrak{c}}(L)$ has received quite a lot of attention in pointfree topology. See [ball18], [joinsofclosed19], [moshier20], [ball19], [ball16]. This is the collection of joins of closed sublocales. Functoriality of the assignment $L\mapsto\mathsf{S}_{\mathfrak{c}}(L)$ is explored in [ball19] for subfit frames, and the authors identify some ways of restricting the objects so that all frame morphisms $f:L\rightarrow M$ lift to frame morphisms. We will use our characterization on how frame maps can be extended to maps of Raney extensions to find conditions on morphisms $f:L\rightarrow M$ between arbitrary frames (not necessarily subfit) to lift, and find that a morphism lifts if and only if for an exact meet $\bigwedge_{i}x_{i}\in L$ we have both that $\bigwedge_{i}f(x_{i})$ is exact and that it equals $f(\bigwedge_{i}x_{i})$.
Another important structure in pointfree topology is the subcoframe $\mathsf{S}_{\mathfrak{o}}(L)\subseteq\mathsf{S}(L)$ of fitted sublocales, that is, intersections of open sublocales. See, for instance, [clementino18], [moshier20]. The structures $\mathsf{S}_{\mathfrak{c}}(L)$ and $\mathsf{S}_{\mathfrak{o}}(L)$ are compared in [moshier20] and in [moshier22]. In this work we show that, for a frame $L$, $\mathsf{S}_{\mathfrak{o}}(L)$ and $\mathsf{S}_{\mathfrak{c}}(L)$ are, respectively, the largest and smallest Raney extensions, showing that these two structures have universal properties which are dual to each other. For $\mathsf{S}_{\mathfrak{o}}(L)$ it is also true that it is the free Raney extension over $L$.
@@
Finally, we will look at canonical extensions of frames, as defined in [jakl20], and show that for a pre-spatial frame its canonical extension is the free Raney extension over it which is algebraic (join-generated by its compact elements), thus showing another universal property of this construction. We will also show that a $T_{0}$ space is sober, resp. strictly sober, if and only if the Raney extension $(\Omega(X),\mathcal{U}(X))$ is compact with respect to the collection of completely prime, resp. Scott-open, filters.
@@ 2.1 Sublocales
    Whenever $s\in S$ and $x\in L$ we have $x\rightarrow s\in S$.
These requirements are equivalent to stating that $S\subseteq L$ is a regular monomorphism in $\mathbf{Loc}$. Observe that the collection of sublocales of a frame is closed under all intersections. We have the following useful fact.
@@
The family $\mathsf{S}(L)$ of all sublocales of $L$ ordered by inclusion is a coframe. Meets in $\mathsf{S}(L)$ are set-theoretical intersections. The top element is $L$ and the bottom element is $\{1\}$. Because $\mathsf{S}(L)$ is a coframe, there is a difference operator on it, dual to Heyting implication, defined for sublocales $S$ and $T$ as $S{\setminus}T=\bigcap\{U\in\mathsf{S}(L):S\subseteq T\cup U\}$. For a sublocale $S$, we denote the element $L{\setminus}S$ as $S^{*}$, and we call it the supplement of $S$.
Open and closed sublocales
We now focus on the pointfree versions of open and closed subspaces. For each $a\in L$, there are an open sublocale and a closed sublocale associated with it. These are, respectively,
$\mathfrak{o}(a)=\{a\rightarrow b:b\in L\},$ $\mathfrak{c}(a)=\uparrow a.$
Open and closed sublocales behave like open and closed subspaces in many respects; we list a few of them below.
@@
    The elements $\mathfrak{o}(a)$ and $\mathfrak{c}(a)$ are complements of each other in $\mathsf{S}(L)$: we have $\mathfrak{o}(a)\cap\mathfrak{c}(a)=$ and $\mathfrak{o}(a)\vee\mathfrak{c}(a)=L$;
@@
The ordered collection of closed sublocales of a frame $L$ is a subcoframe of $\mathsf{S}(L)$. In fact, this coframe is anti-isomorphic to $L$. The collection of joins of closed sublocales of a frame is a frame and is called $\mathsf{S}_{\mathfrak{c}}(L)$. Open sublocales form a subframe of $\mathsf{S}(L)$, and this is isomorphic to $L$. The sublocales which are intersections of open sublocales are called fitted. These form a subcolocale of $\mathsf{S}(L)$ which is called $\mathsf{S}_{\mathfrak{o}}(L)$.
@@ 2.2 Saturated sets and fitted sublocales
[…]
$F\mapsto\bigcap\{\mathfrak{o}(f):f\in F\}.$
This maximally restricts to an anti-isomorphism between fitted sublocales and strongly exact filters (see [moshier20]. Recall that a meet $\bigwedge_{i}x_{i}$ is strongly exact if for all $y\in L$ we have that $x_{i}\rightarrow y=y$ implies $(\bigwedge_{i}x_{i})\rightarrow y=y$, and that a filter is strongly exact if it is closed under strongly exact meets. We call $\mathsf{Filt}_{\mathcal{SE}}(L)$ the ordered collection of strongly exact filters. This is a frame where meets are computed as intersections, and additionally it is a sublocale of $\mathsf{Filt}(L)$.
Lemma 2.5.
For a filter $F\subseteq L$, the following are equivalent.
  1. 1.
    $F$ is strongly exact.
  2. 2.
    $\bigcap_{f\in F}\mathfrak{o}(f)\subseteq\mathfrak{o}(x)$ is equivalent to $x\in F$.
  3. 3.
[…]
  4. 4.
    $F=\{x\in L:S\subseteq\mathfrak{o}(x)\}$ for some sublocale $S\subseteq L$.
[…]
@@
We have an isomorphism of coframes
@@
In this paragraph, we refer to [jakl24], and mention the main results from that we are going to use. We will refer to several important concrete collections of filters. Since the collection $\mathsf{Filt}(L)$ is a frame, there is a Heyting operation $\rightarrow$ on it. In the following, whenever we write $F\rightarrow G$ for two filters $F$ and $G$, it will be understood that we are referring to this operation. Notice that for a frame $L$ and for $a,b\in L$ we have
@@
In particular, note that this means that for any $a\in L$ we have $\neg{\uparrow}a={\uparrow}a\to\{1\}=\{x\in L:x\vee a=1\}$. We say that a filter is regular if it is a regular element in the frame of filters (that is, if it is of the form $\neg F$ for some filter $F$). We call $\mathsf{Filt}_{\mathcal{R}}(L)$ the ordered collection of regular filters. Note that $\mathsf{Filt}_{\mathcal{R}}(L)\subseteq\mathsf{Filt}(L)$ is the Booleanization of the frame of $\mathsf{Filt}(L)$. Regular filters, too, have a useful concrete characterization.
@@
For any frame $L$, we have the following poset of sublocale inclusions:
@@
Finally, we have the following characterizations of frame properties in terms of collections of filters.

### v2

@@ Abstract
We introduce a pointfree version of Raney duality. Our objects are Raney extensions of frames, pairs $(L,C)$ where $C$ is a coframe and $L\subseteq C$ is a subframe that meet-generates it and whose embedding preserves strongly exact meets. We show that there is a dual adjunction between $\mathbf{Raney}$ and $\mathbf{Top}$, with all $T_{0}$ spaces as fixpoints, assigning to a space $X$ the pair $(\Omega(X),\mathcal{U}(X))$, with $\mathcal{U}(X)$ are the intersections of open sets. We show that for every Raney extension $(L,C)$ there are subcolocale inclusions $\mathsf{S}_{\mathfrak{c}}(L)^{op}\subseteq C\subseteq\mathsf{S}_{\mathfrak{o}}(L)$ where $\mathsf{S}_{\mathfrak{o}}(L)$ is the coframe of fitted sublocales and $\mathsf{S}_{\mathfrak{c}}(L)$ is the frame of joins of closed sublocales. We thus exhibit a symmetry between these two well-studied structures in pointfree topology. The spectra of these are, respectively, the classical spectrum $\mathsf{pt}(L)$ of the underlying frame and its $T_{D}$ spectrum $\mathsf{pt}_{D}(L)$. This confirms the view advanced in [banaschewskitd] that sobriety and the $T_{D}$ property are mirror images of each other, and suggests that the symmetry above is a pointfree view of it. All Raney extensions satisfy some variation of the properties density and compactness from the theory of canonical extensions. We characterize sobriety, the $T_{1}$, and the $T_{D}$ axioms in terms of density and compactness of $(\Omega(X),\mathcal{U}(X))$. We characterize frame morphisms $f:L\to M$ that extend to Raney morphisms $\overline{f}:(L,C)\to(M,D)$. We thus obtain a characterization of morphisms of frames $f:L\to M$ which extend to frame morphisms $\overline{f}:\mathsf{S}_{\mathfrak{c}}(L)\to\mathsf{S}_{\mathfrak{c}}(M)$, answering a question posed in [ball14b]. We show the existence of the free Raney extension over a frame. We show that all Raney extensions admit a sober coreflection. Restricting morphisms of $\mathbf{Raney}$ to exact morphisms gives both cofree objects and $T_{D}$ reflections. Finally, we show that the canonical extension of a locally compact frame (introduced in [jakl20]) is the free algebraic Raney extension. We also give a new view of $T_{D}$ duality: in contrast with the frame case, $T_{D}$ spaces are are full subcategory of $\mathbf{Raney}$, with no need to restrict morphisms.
@@ 1 Introduction
In this work, we study an algebraic, pointfree version of the embedding of a frame of open sets into a lattice of saturated sets. Our constructions are inspired by Raney duality, as illustrated in [bezhanishvili20]. In Raney duality, we map a space $X$ to the pair $(\Omega(X),\mathcal{U}(X))$ where $\mathcal{U}(X)$ is the lattice of saturated sets [^1] Saturated sets are intersections of open sets.. In this duality, on the algebraic side all objects of the category have, so to speak, enough points. Every Raney algebra is of the form $(\Omega(X),\mathcal{U}(X))$ for some space $X$. In this work, we extend the category of Raney algebras to include pointfree objects. We do so by taking as objects generalizations of the canonical extension construction for distributive lattices. Canonical extensions for Boolean algebras were introduced by Jónsson and Tarski (see [jonsson51i] and [jonsson51ii]) in dealing with Boolean algebras with operators. They have later proven to be useful in theoretical Computer Science and Logic, and have been generalized to various settings. Canonical extension may be seen as an algebraic and pointfree version of the lattice of compact opens into the lattice of saturated sets of a Stone space. Canonical extensions have also been introduced for distributive lattices. On this topic, we refer the reader to [gehrke94], [gehrke01], and [gehrke08]. For distributive lattices this represents the embedding of the lattice of compact open sets of a coherent space into the lattice of saturated sets. For a distributive lattice, its canonical extension is unique. In [jakl20], canonical extension is introduced for locally compact frames. For general frames, we claim, there is no unique way of extending a frame $L$ to a pointfree lattice of saturated sets. For a concrete space $X$, the structure of $\mathcal{U}(X)$ will depend, for example, on how many of the points of $\Omega(X)$ are realized as concrete points of $X$.
@@
In Section 3 we introduce the main structures. We will consider as pointfree spaces Raney extensions, pairs $(L,C)$ where $C$ is a coframe and $L\subseteq C$ is a frame which meet-generates $C$ and such that the embedding preserves the frame operations together with strongly exact meets [^2] Strongly exact meets may be seen as the pointfree version of those intersections of open sets which are open, as they are characterized as those that give collections of open sublocales with open intersections. Because a meet of a collection $\{U_{i}:i\in I\}$ of opens in general is calculated as the interior of $\bigcap_{i}U_{i}$, these are exactly the meets that are preserved by the embedding $\Omega(X)\subseteq\mathcal{U}(X)$. . Every Raney extension will satisfy a generalization of the property density and compactness from the theory of canonical extensions. We build on the work in [jakl24] connecting canonical extensions and pointfree topology. In pointfree topology, a structure of primary importance is the structure $\mathsf{S}(L)$ of sublocales (pointfree subspaces) of a frame $L$. The structures $\mathsf{S}_{\mathfrak{c}}(L)$ of joins of closed sublocales and $\mathsf{S}_{\mathfrak{o}}(L)$ of fitted sublocales have been widely studied, and these are compared in [moshier20] and in [moshier22]. We will show that these are the largest and smallest Raney extensions on a frame $L$, respectively.
In Section 4 we will show that there is an adjunction $\Omega_{R}:\mathbf{Top}\leftrightarrows\mathbf{Raney}^{op}:\mathsf{pt}_{R}$ extending Raney duality. As for canonical extensions, the points of a Raney extension $(L,C)$ are completely join-prime elements of $C$. We compute the spectrum of the largest and smallest Raney extensions $(L,\mathsf{S}_{\mathfrak{o}}(L))$ and $(L,\mathsf{S}_{\mathfrak{c}}(L)^{op})$, and discover that these are, respectively, the spectrum $\mathsf{pt}(L)$ of the underlying frame and its $T_{D}$ spectrum $\mathsf{pt}_{D}(L)$. The $T_{D}$ spectrum was introduced in [banaschewskitd], where a duality is shown between $T_{D}$ spaces and the category $\mathbf{Frm}_{D}$, obtained by restricting morphisms in $\mathbf{Frm}$. In our case, $T_{D}$ spaces are faithfully represented in $\mathbf{Raney}$ as the subcategory of the (spatial) Raney extensions of the form $(L,\mathsf{S}_{\mathfrak{c}}(L)^{op})$, with no need to restrict morphisms.
In Section 5 we exploit the duality to characterize topological properties such as sobriety, the $T_{D}$ and the $T_{1}$ axioms in terms of density and compactness of their Raney extensions. In pointfree topology, the $T_{1}$ axiom has several different translations; weakest is subfitness, see for example Chapter V of [picadopultr2012frames] or Chapter II of [picado21]. A Raney extension $(L,C)$ is defined to be $T_{1}$ if $C$ is Boolean, just like a space $X$ is $T_{1}$ if and only if $\mathcal{U}(X)$ is the powerset. We connect the two views by showing that a frame is subfit if and only if it admits a $T_{1}$ Raney extension. We thus exhibit subfitness as the weakest possible frame version of the $T_{1}$ axiom. The notions of sober and $T_{D}$ objects have no counterpart in the frame setting. Another axiom which has been studied quite extensively is scatteredness. Scatteredness for a frame $L$ is defined in [plewe00] and [plewe02] as the property that $\mathsf{S}(L)$ is Boolean. In [ball16], the authors characterize the frames for which $\mathsf{S}_{\mathfrak{c}}(L)=\mathsf{S}(L)$ as those subfit frames such that they are scattered. Here, we show that subfit frames which are scattered coincide with those subfit frames with unique Raney extensions.
In Section 6, we will characterize frame morphisms $f:L\to M$ that extend to Raney morphisms $\overline{f}:(L,C)\to(M,D)$. In this case, the requirement is weaker than in the theory of canonical extension, where extensions of morphisms are required to preserves all meets and all joins. We uswe this to answer the question posed in [ball19] of what morphisms of frames $f:L\to M$ lift to frame morphisms $\mathsf{S}_{\mathfrak{c}}(f):\mathsf{S}_{\mathfrak{c}}(L)\to\mathsf{S}_{\mathfrak{c}}(M)$. We call them exact frame morphisms. We also use the result to show that every frame admits a free Raney extensions on it. Restricting morphisms on $\mathbf{Raney}$ suitably also gives $T_{D}$ reflections. We also show that every Raney extension has a sober coreflection. Both these results have no counterpart in $\mathbf{Frm}$.
@@
In Section 7, finally, we look at two special topics. We will look at canonical extensions of frames, as defined in [jakl20], and show that for a pre-spatial frame its canonical extension is the free Raney extension over it which is algebraic (join-generated by its compact elements). We will also show that a $T_{0}$ space is sober, resp. strictly sober, if and only if the Raney extension $(\Omega(X),\mathcal{U}(X))$ is compact with respect to the collection of completely prime, resp. Scott-open, filters. We also look at the category of frames with exact morphisms $\mathbf{Frm}_{\mathcal{E}}$. We show that this is isomorphic to the subcategory $\mathbf{Raney_{D}}$ of $\mathbf{Raney}$, defined as $\{(L,\mathsf{S}_{\mathfrak{c}}(L)^{op}):L\in\mathbf{Frm}\}$. This gives a new view of $T_{D}$ duality, pictured below. Note that $\mathsf{S}_{\mathfrak{c}}:\mathbf{Frm}_{\mathcal{E}}\to\mathbf{Raney_{D}}$ is an equivalence.
${\mathbf{Frm}_{\mathcal{E}}}$${\mathbf{Top}_{D}}$${\mathbf{Raney_{D}},}$$\scriptstyle{\mathsf{pt}_{D}}$$\scriptstyle{\mathsf{S}_{\mathfrak{c}}}$$\scriptstyle{\mathsf{pt}_{R}}$
@@ 2.1 Sublocales
    Whenever $s\in S$ and $x\in L$, $x\rightarrow s\in S$.
These requirements are equivalent to stating that $S\subseteq L$ is a regular monomorphism in $\mathbf{Loc}$. Observe that the collection of sublocales of a frame is closed under all intersections. The following is a useful fact.
@@
The family $\mathsf{S}(L)$ of all sublocales of $L$ ordered by inclusion is a coframe. Meets in $\mathsf{S}(L)$ are set-theoretical intersections. For a subset $X\subseteq L$, we denote as $\mathcal{S}(X)$ the smallest sublocale containing $X$. In the following, $\mathcal{M}(-)$ denotes closure under meets.
Lemma 2.2.
For a frame $L$ and for $X\subseteq L$, $\mathcal{S}(X)=\mathcal{M}(\{a\rightarrow x:a\in L,x\in X\})$.
The top element is $L$ and the bottom element is $\{1\}$. Because $\mathsf{S}(L)$ is a coframe, there is a difference operator on it, dual to Heyting implication, defined for sublocales $S$ and $T$ as $S{\setminus}T=\bigcap\{U\in\mathsf{S}(L):S\subseteq T\cup U\}$. For a sublocale $S$, we denote the element $L{\setminus}S$ as $S^{*}$, and we call it the supplement of $S$. For each $a\in L$, there are an open sublocale and a closed sublocale associated with it. These are, respectively, $\mathfrak{o}(a)=\{a\rightarrow b:b\in L\}$ and $\mathfrak{c}(a)=\uparrow a$. We will need a few facts about open and closed sublocales, which we gather here.
@@
    The elements $\mathfrak{o}(a)$ and $\mathfrak{c}(a)$ are complements of each other in $\mathsf{S}(L)$;
@@
Every sublocale can be written as an intersection of sublocales of the form $\mathfrak{o}(x)\vee\mathfrak{c}(y)$. A sublocale is fitted if it is an intersection of open sublocales. We call $\mathsf{S}_{\mathfrak{o}}(L)$ the ordered collection of all fitted sublocales. We call $\mathsf{S}_{\mathfrak{c}}(L)$ the ordered collection of joins of closed sublocales. This collection is studied in [joinsofclosed19]. For a coframe $C$, we say that an element $c\in C$ is linear if $\bigvee_{i}(x_{i}\wedge c)=\bigvee_{i}x_{i}\wedge c$ for any collection $x_{i}\in C$.
Lemma 2.4.
Complemented elements of a coframe are linear. In particular, in $\mathsf{S}(L)$ open and closed sublocales are linear.
Also particularly important are Boolean sublocales. For an element $a\in L$ the sublocale $\{x\to a:x\in L\}$, denoted as $\mathfrak{b}(a)$, is the smallest sublocale containing $a$. A sublocale is a Boolean algebra if and only if it is of this form for some $a\in L$. An element $p\in L$ is prime when $x\wedge y\leq p$ implies either $x\leq p$ or $y\leq p$, for all $x,y\in L$. Elements of the form $\mathfrak{b}(p)$ are also called two-element sublocales, as for $p$ prime $\mathfrak{b}(p)=\{1,p\}$.
Lemma 2.5.
For a frame $L$, the following hold for each prime $p\in L$ and all elements $x,y\in L$.
  • $x\to p=1$ if $x\leq p$, and $x\to p=p$ if $x\nleq p$.
  • $\mathfrak{b}(p)\subseteq\mathfrak{o}(x)$ if and only if $x\nleq p$.
  • The element $\mathfrak{b}(p)$ is completely join-prime in $\mathsf{S}(L)$.
  • The prime elements of $\mathsf{S}(L)$ are the sublocales of the form $\mathfrak{b}(p)$ for some prime $p\in L$.
@@ 2.2 Saturated sets and fitted sublocales
There is an isomorphism of coframes
@@
The following results give topological intuition on exactness and strong exactness.
Proposition 2.9.
For every space $X$, strongly exact meets in $\Omega(X)$ are open sets. This means that the embedding $\Omega(X)\subseteq\mathcal{U}(X)$ preserves strongly exact meets.
Theorem 2.10 (see [ball14], Theorem 5.2.3).
A $T_{0}$ space is $T_{D}$ if and only if for every exact meet $\bigwedge_{i}U_{i}$ in $\Omega(X)$ this equals $\bigcap_{i}U_{i}$. This is equivalent to the embedding $\Omega(X)\subseteq\mathcal{U}(X)$ preserving exact meets.
@@
In this paragraph, we refer to [jakl24], and mention the main results from that we are going to use. We will refer to several important concrete collections of filters. Since the collection $\mathsf{Filt}(L)$ is a frame, there is a Heyting operation $\rightarrow$ on it. Notice that for a frame $L$ and for $a,b\in L$
@@
In particular, note that this means that, for any $a\in L$, $\neg{\uparrow}a={\uparrow}a\to\{1\}=\{x\in L:x\vee a=1\}$. We say that a filter is regular if it is a regular element in the frame of filters (that is, if it is of the form $\neg F$ for some filter $F$). We call $\mathsf{Filt}_{\mathcal{R}}(L)$ the ordered collection of regular filters. Note that $\mathsf{Filt}_{\mathcal{R}}(L)\subseteq\mathsf{Filt}(L)$ is the Booleanization of the frame of $\mathsf{Filt}(L)$. Regular filters, too, have a useful concrete characterization.
@@
For any frame $L$, there is the following poset of sublocale inclusions:
@@
Finally, the following characterizations of frame properties in terms of collections of filters.